There are six pairs of integers that may be chosen from the four positive integers a < b < c < d. For each such pair, the positive difference between the two integers is calculated. Prove that the product of these six positive differences is divisible by 12.
Answer by ikleyn(53763)

.

Consider these integers moduo 4.



If they are all different modulo 4, then they are 

    p= 0 (mod4),  q= 1 (mod4), r= 2(mod4) and s= 3(mod4).



In this case, we can choose two pairs from 6 differences, that produce 2 (mod4):  r-p and s-q.



If, in opposite, they are not all different (mod4), then we have at least one pair producing difference 0(mod4).



So, in any case, we have at least one multiplier 0 (mod4).



Similar logic works for (mod3).



Of the four remainders mod3, we have at least one pair with the difference 0 (mod3).



Thus, the statement is proved.